I have tried to solve this for several days, searching everywhere for help. This is my last try, please HELP me! Here is the scenario:
Select any two integers between -12 and +12 which will become solutions to a system of two equations.
Write two equations that have your two integers as solutions. Show how you built the equations using your integers. Your solution and equations should not be the same as those of other students or the textbook. There are infinite possibilities.
Solve your system of equations by the addition/subtraction method. Make sure you show the necessary 5 steps.

Answer by Theo(13342)

let x = 5
let y = 6
your equations will be of the form ax + by = c
you can pick any a and any b and solve for c using the values of x and y that you chose.
i chose:
2x + 6y = 46
7x - 3y = 17
since x = 5 and y = 6, 2x + 6y is equal to 2(5) + 6(6) which is equal to 10 + 36 which is equal to 46.
i did the same for the second equation.
the 2 equations are:
2x + 6y = 46
7x - 3y = 17
to solve these by elimination method (adding or subtracting equations from each other), you need to get the x or the y variables to cancel out.
multiplying the second equation by 2 will allow the y variable to cancel out.
2 times the second equation gets you 7x - 3y = 17 to become 14x - 6y = 34
your 2 equation to now solve are:
2x + 6y = 46
14x - 6y = 34
add these 2 equations together to get:
16x = 80
divide both sides by 16 to get x = 5
now you can substitute for x in one of the equations and solve for y to get y = 6.